Grade 12Physical Chemistry

11.1g. of CaCl2 is present in 100ml of the aqueous solution. The chloride ion concentration is

shiva kumar
7 Years agoGrade 12

2 Answers

aparazita sharma
7 Years ago
Si
aparazita sharma
7 Years ago
since in ques we have 11.1 g caclwe will first calculate no of moles corresponding to 11.1 g ehich is equal to 0.1 mole as no of moles = given mass/ molar mass 
and molar mass = 40+2*35.5 =111g 
so no of moles = 11.1 / 111 = 0.1 mole 
now as CaCl2 is in aq medium it will dissociate as 
            CaCl2 (aq)   gives    Ca2+(aq) +  2 Cl-(aq)
now since 1 mole is giving 2 mole cl- therefore 0.1 mole cacl2 will give 0.2 mole cl- ion   . We can calculate concentration as Molarity = no of moles / vol of solution in L    therefore conc of cl- ion will be 0.2 / 0.1 = 2M (as 100 ml  = 0.1 L) (100 / 1000)